Please tell me how to calculate no.of squares in a chess board having common side ? 

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
we need to count how many neighboring squares have a common side.
1. We have 4 corner squares with 2 neighbors each
2. we have 4 sides with 6 squares on each side with 3 neighbors each
3. we have 6*6 squares in the center with 4 neighbors
=(4*2 + 4*6*3 + 6*6*4)=8+72+144=224